# Impact of Reactive Jamming Attacks on LoRaWAN: a Theoretical and Experimental Study

### v1

@@ V Experimental assessment
The GNU Radio implementation developed for the experiment is hosted on GitHub at [20] for research purposes only, with the important disclaimer that any form of jamming is illegal. The reactive jammer is divided into three complementary blocks:
(i) FrameDetect block: This block is responsible for detecting the beginning of a LoRa transmission, using physical layer parameters such as the central frequency, bandwidth, and SF. Upon detecting a frame, FrameDetect attempts to recover its header and checks the header’s CRC for validity. If the header is valid, it sends a message containing the SF and CR to the trigger block.
(ii) Trigger block: This block determines the number of jamming symbols to generate for each CR value. When it receives a message from the FrameDetect block, it randomly generates symbol values according to the prior configuration and sends them to the jammer block.
(iii) Jammer block: This block modulates the received symbol values according to its physical layer configuration (SF, bandwidth, sampling rate) and transmits the modulated symbols to the next block, which could be an RF transmitter.
To address the synchronization issue identified in [14], a symbol repetition method is used. If the configured number of jamming symbols exceeds one, the trigger block generates the first symbol value randomly and repeats it for the required number of times.
The experimental setup is shown in Fig. 7. The transmitting end node consists of a USRP-2920 connected to a Raspberry Pi 4 running the LoRa implementation from [16]. The jammer is made up of a USRP-2920 connected to a 3 GHz, 6-core laptop running our jammer implementation, and the receiver is a TTGO-LoRa32.
The end node transmits at 10 dBm, while the jammer is configured with a transmission power of 13 dBm. The experiment targets SF values of 7 and 10, and 1000 frames are transmitted for each combination of CR and the number of jamming symbols. The resulting frame success rates are displayed in Figs. 8a and 8b, along with their theoretical counterparts. The experimental frame success rate’s lower bound is 0 frames received out of the 1000, but for plotting purposes, we set this lower bound to $10^{-5}$ on the logarithmic scale.
@@
Additionally, the experimental curves closely follow the theoretical ones, with a slight deviation for CR 3&4 and 1 jamming symbol. This difference arises from the initial assumption that the jamming symbol is perfectly aligned with the legitimate symbol, which is not the case in practice. Instead, the jamming symbol is split and affects two successive legitimate symbols, potentially corrupting more than one bit per codeword.
@@
Figure 8: Experimental Frame success rate under jamming: (a) SF7, (b) SF10
@@
Finally, Tables 1 and 2 demonstrate the effectiveness of our jamming approach. Table 1 presents the signal-to-jamming ratio (SJR) for each combination of spreading factor (SF) and coding rate (CR). It is calculated as the ratio of the number of legitimate symbols, weighted by transmission power, to the number of jamming symbols, weighted by jamming power. The legitimate symbols correspond to physical payloads of 180 bytes for SF7 and 60 bytes for SF10, both of which are below the maximum allowable payload size. Table 2, on the other hand, shows the gain in jamming symbols achieved by our method compared to a conventional LoRa transceiver, which transmits 18.25 symbols without a payload. The computed gains would be even greater if the payload were included for the conventional case. This once again highlights how susceptible the standard LoRa modulation is when faced with a well-designed jammer.
@@ VI Conclusion
In this paper, we explored reactive jamming attacks against LoRaWAN, focusing on minimizing jammer exposure time. We identified and analyzed the key components of the LoRa transceiver chain that mitigate jamming attacks. Our analysis demonstrates how the frame success rate is impacted by the number of jamming symbols transmitted. It also shows that transmitting fewer than one interleaving block is enough to significantly reduce the frame success rate, provided the jamming power is sufficient for the jammer’s symbol to be misinterpreted as the legitimate symbol by the receiver. Additionally, we presented a GNU Radio implementation of the jammer to validate our analysis. The deterministic and known preamble pattern of LoRa makes reactive jamming feasible and straightforward. Future work will focus on developing a stealthy preamble structure for LoRa as a countermeasure to reactive jamming.

### v2

@@ V Experimental Assessment
The GNU Radio implementation developed for the experiment is hosted on GitHub at [19] for research purposes only, with the important disclaimer that any form of jamming is illegal. The reactive jammer is divided into three complementary blocks: (i) The FrameDetect block detects a LoRa transmission, based on the central frequency, bandwidth, and SF. Upon detection, it attempts to recover its header and extract the CR. If the header is valid, it sends a message containing the SF and CR to the trigger block. (ii) The Trigger block determines the number of jamming symbols to generate for each CR value. When it receives a message from the FrameDetect block, it randomly generates a symbol value, duplicates it – repeated-jamming – according to the prior configuration, and sends it to the jammer block. (iii) The Jammer block modulates the received symbol values according to its physical layer configuration (SF, bandwidth, sampling rate) and transmits the modulated symbols to the next block, which could be an RF transmitter.
The experimental setup is shown in Fig. 7. The transmitting end node is a USRP-2920 connected to a Raspberry Pi 4 running the LoRa implementation from [20]. The jammer is a USRP-2920 connected to a laptop running our jammer implementation, and the receiver is a TTGO-LoRa32.
The end node transmits at 10 dBm, while the jammer is configured with a transmission power of 13 dBm. The experiment considers SF 7 and SF 10, and 1000 frames are transmitted for each combination of CR and number of jamming symbols. The resulting frame success rates are displayed in Figs. 8a and 8b, along with their theoretical repeated-symbol jamming counterparts. For cases where none out of the 1000 frames were received, we set the success rate to $10^{-5}$ for better illustration.
@@
Additionally, the experimental curves closely follow the theoretical ones, with a slight deviation for CR 3&4 and 1 jamming symbol. This difference arises from the initial assumption made on the jamming symbol offset distribution. It therefore remains an interesting task to study the real distribution of that offset.
@@
Figure 8: Experimental Frame success rate under jamming $V_{A}=3\mbox{ dB}$: (a) SF7, (b) SF10
@@
Finally, Tables I and II demonstrate the effectiveness of our jamming approach. Table I presents the signal-to-jamming ratio (SJR) for each combination of spreading factor (SF) and coding rate (CR). It is calculated as the ratio of the number of legitimate symbols, weighted by transmission power, to the number of jamming symbols, weighted by jamming power. The legitimate symbols correspond to physical payloads of 180 bytes for SF7 and 60 bytes for SF10, both of which are below the maximum allowable payload size. Table II, on the other hand, shows the gain in jamming symbols achieved by our method compared to a jammer based on a conventional LoRa transceiver, which transmits 18.25 symbols without a payload. This once again highlights how susceptible the standard LoRa modulation is when faced with a well-designed jammer.
@@ VI Conclusion
In this paper, we explored reactive jamming attacks against LoRaWAN, focusing on minimizing jammer exposure time. Our analysis demonstrates how the frame success rate is impacted by the number of jamming symbols in the case of synchronized jamming, non-synchronized jamming, and repeated-symbol jamming. The results show that transmitting fewer than one interleaving block can be enough to significantly reduce the frame success rate, provided the jamming-to-signal magnitude ratio is sufficient for the jammer’s symbol to be misinterpreted as the legitimate symbol by the receiver. Additionally, we presented a GNU Radio implementation of the jammer to validate our analysis. The deterministic and known preamble pattern of LoRa makes reactive jamming feasible and straightforward. Future work will focus on developing a stealthy preamble structure for LoRa as a countermeasure to reactive jamming.